A force vector points at an angle of 54.4° above the +x axis. It has a y component of +280 newtons. Calculate the magnitude of the force vector. Submit Answer Tries 0/99
Answer #1

Let a fore F is making an agle 54.5° above the x axis shows in figure..グ Fsins4s个 S A

X - component = Fcos(54.5°)....Eqn(i)

Y - component = Fsin(54.5°).....Eqn(ii)

Now in question Y comppnent is given as 280N

so , From equation (ii)

280 = Fsin(54.5°)

solving for F , F = 280/(sin54.5°) = 343.93 N

so now we can calculate its X - Component

X = Fcos(54.5° ) = 343.93cos(54.5°) = 199.722 N

Hence magnitude |F| = √(X^2 + Y^2) = 397.71N...Answer.

Please thumbs up?.